# Exact Subgraph Isomorphism Network for Predictive Graph Mining

### v1

@@ Abstract
In the graph-level prediction task (predict a label for a given graph), the information contained in subgraphs of the input graph plays a key role. In this paper, we propose Exact subgraph Isomorphism Network (EIN), which combines the exact subgraph enumeration, a neural network, and a sparse regularization. In general, building a graph-level prediction model achieving high discriminative ability along with interpretability is still a challenging problem. Our combination of the subgraph enumeration and neural network contributes to high discriminative ability about the subgraph structure of the input graph. Further, the sparse regularization in EIN enables us 1) to derive an effective pruning strategy that mitigates computational difficulty of the enumeration while maintaining the prediction performance, and 2) to identify important subgraphs that contributes to high interpretability. We empirically show that EIN has sufficiently high prediction performance compared with standard graph neural network models, and also, we show examples of post-hoc analysis based on the selected subgraphs.
@@ 1 Introduction
EIN consists of Graph Mining Layer (GML) and Feed Forward Network (FFN). GML implements a mechanism to select only a small number of subgraphs necessary for the prediction, achieved through the group-sparse regularization. The group sparse regularization (Yuan and Lin,, 2006) is a well-known regularizer for the group-wise selection of variables. We regard a set of neural network weights from an input unit (corresponding to one subgraph) to the next layer as a group, by which the adaptive subgraph selection can be realized. Only the selected subgraphs are used in the subsequent FFN that learns the dependency between the selected subgraphs and the prediction target. Unfortunately, although the number of subgraphs becomes small in the trained model, this approach still suffers from the computational difficulty because all the candidate subgraphs are still required to consider during the optimization process.
To tackle the computational difficulty, we combine a subgraph enumeration by graph mining (Yan and Han,, 2002) and the proximal gradient optimization (Teboulle,, 2017; Beck and Teboulle,, 2009), by which a pruning strategy for unnecessarily subgraphs can be derived, shown in 1 (c). The proximal gradient is a standard approach for the sparse modeling, in which parameters are updated through a proximal projection that typically results in a thresholding operation. This thresholding operation reveals that if the norm of the gradient corresponding to a subgraph $H$ is less than certain threshold, parameters for that subgraph is not required to update. We show that, by deriving an upper bound of the norm of the gradient, an efficient pruning strategy for the subgraph enumeration can be constructed. Our pruning strategy has the following two important implications. First, this pruning enables us to train EIN without enumerating all the candidate subgraphs, which makes EIN computationally tractable. Second, our pruning strategy maintains the quality of the prediction compared with when we do not perform the pruning. This is because we only omit the computations that does not have any effect on the prediction.
@@
  • We propose EIN, which is a neural network model that uses the exact subgraph isomorphism feature. Through a group sparse regularization, we formulate EIN so that a small number of subgraphs can be identified by which an insight about the important substructure can be extracted for the given graph-level prediction task.
  • We show that the combination of the subgraph enumeration and the proximal gradient with backpropagation can derive an efficient pruning strategy that makes the EIN training computationally tractable. We further reveal that our pruning strategy does not degrade the prediction quality.
@@ 2 Proposed Method: Exact Subgraph Isomorphism Network
Our proposed method, called Exact subgraph Isomorphism Network (EIN), considers the classification problem of a graph $G\in{\cal G}$, where $\cal G$} is a set of labeled graphs. $G$ consists of a set of nodes and edges between nodes and each node can have a categorical label. The training data is $\{(G_{i},y_{i})\}_{i\in[n]}$, where $y_{i}\in{\cal Y}$ is a graph label, $n$ is the number of instances, and $n]=\{1,\ldots,n\}$. Here, although we only focus on the classification problem, the regression problem can also be handled by just replacing the loss function.
@@ 2.1 Model Definition
As shown in Fig. 1 (b), EIN consists of Graph Mining Layer (GML) and Feed Forward Network (FFN). Let $K$ be the number of output units of GML. We define $\mathrm{GML}:{\cal G}\rightarrow\mathbb{R}^{K}$ as follows.
@@
$\boldsymbol{h}$ $=\sum_{H\in{\cal H}}\boldsymbol{\beta}_{H}\psi_{H}(G_{i})+\boldsymbol{b},$
where $\boldsymbol{B}=[\boldsymbol{\beta}_{1},\dots,\boldsymbol{\beta}_{|{\cal H}|}]\in\mathbb{R}^{K\times|{\cal H}|}$ and $\boldsymbol{b}\in\mathbb{R}^{K}$ are parameters, and $\sigma:\mathbb{R}^{K}\rightarrow\mathbb{R}^{K}$ is an activation function. Each $\boldsymbol{\beta}_{H}$ can be seen as a representation corresponding to each subgraph $H$. The entire prediction model is defined as
@@
where $\boldsymbol{\Theta}$ is parameters of FFN. We optimize the parameters $\boldsymbol{B},\boldsymbol{b}$, and $\boldsymbol{\Theta}$ through the following optimization problem in which the group sparse penalty (Yuan and Lin,, 2006) is imposed on $\boldsymbol{\beta}_{H}$:
$\min_{\boldsymbol{B},\boldsymbol{b},\boldsymbol{\Theta}}\ \sum_{i=1}^{n}\ell(y_{i},f(G_{i}))+\lambda\sum_{H\in{\cal H}}\|\boldsymbol{\beta}_{H}\|_{2}$ (2)
where $\ell$ is a differentiable loss function and $\lambda$ is a regularization parameter. We here use the cross-entropy loss for $\ell$. The group-wise penalty $\|\boldsymbol{\beta}_{H}\|_{2}$ results in $\boldsymbol{\beta}_{H}=\boldsymbol{0}$ for many unnecessarily $H$ at the solution of (2), by which we can identify important predictive subgraphs as $\{H\mid\boldsymbol{\beta}_{H}\neq\boldsymbol{0}\}$. However, since the size of $\boldsymbol{B}$, i.e., $K\times|{\cal H}|$, is quite large, naïve optimization of (2) can be difficult.
@@ 2.2 Optimization
Our optimization algorithm is based on the block coordinate descent (Xu and Yin,, 2017), in which each one of $\boldsymbol{B},\boldsymbol{b}$, and $\boldsymbol{\Theta}$ are updated alternately while the other two parameters are fixed. Since $\boldsymbol{B}$ has the group sparse penalty, we apply the well-known proximal gradient method (Beck and Teboulle,, 2009; Parikh and Boyd,, 2014). On the other hand, $\boldsymbol{b}$ and $\boldsymbol{\Theta}$ have no sparse penalty and we can simply apply the standard gradient descent.
Let $\boldsymbol{g}_{H}={\frac{\partial}{\partial\boldsymbol{\beta}_{H}}}\sum_{i=1}^{n}\ell(y_{i},f(G_{i}))$ be the gradient of the loss function with respect to $\boldsymbol{\beta}_{H}$ and
[…]
be the proximal projection with respect to $\eta\lambda\|\cdot\|_{2}$ for a given vector $\boldsymbol{a}$ with the step size $\eta$. Then, the update of each parameter is defined as
$\boldsymbol{\beta}^{\rm(new)}_{H}$ $\leftarrow\mathrm{prox}\left(\boldsymbol{\beta}_{H}-\eta\ \boldsymbol{g}_{H}\right)\text{ for }H\in{\cal H},$ (4)
$\boldsymbol{b}^{\rm(new)}$ $\leftarrow\boldsymbol{b}-\alpha\sum_{i=1}^{n}{\frac{\partial\ell(y_{i},f(G_{i}))}{\partial\boldsymbol{b}}},$ (5)
$\boldsymbol{\Theta}^{\rm(new)}$ $\leftarrow\boldsymbol{\Theta}-\gamma\sum_{i=1}^{n}{\frac{\partial\ell(y_{i},f(G_{i}))}{\partial\boldsymbol{\Theta}}},$ (6)
[…]
@@ 2.2.1 Backpropagation with Gradient Pruning
[…]
$\mathrm{prox}\left(\boldsymbol{\beta}_{H}-\eta\ \boldsymbol{g}_{H}\right)=\boldsymbol{0}\ \text{ for }\ H\in\{H\mid\|\boldsymbol{g}_{H}\|_{2}\leq\lambda\text{ and }H\in\overline{{\cal W}}\}$ (7)
This means that if $\|\boldsymbol{g}_{H}\|_{2}\leq\lambda\text{ and }H\in\overline{{\cal W}}$ are satisfied, both of the current $\boldsymbol{\beta}_{H}$ and the updated $\boldsymbol{\beta}_{H}^{\rm new}$ are zero. From this observation, we incrementally update $\cal W$} as
${\cal W}\leftarrow{\cal W}\cup\left\{H\ \middle|\ \|\boldsymbol{g}_{H}\|_{2}>\lambda,\forall H\in\overline{{\cal W}}\right\},$ (8)
and perform the update (4) only for $H\in{\cal W}$. However, evaluating $\|\boldsymbol{g}_{H}\|_{2}>\lambda$ for all $H\in\overline{{\cal W}}$ is computationally demanding. To avoid this difficulty, the following theorem plays a key role:
@@
Let $H^{\prime}\sqsupseteq H$ for $H,H^{\prime}\in\overline{{\cal W}}$, and $\delta_{ik}={\frac{\partial\ell(y_{i},f(G_{i}))}{\partial h_{k}}}$. Then, we have
$\|\boldsymbol{g}_{H^{\prime}}\|_{2}\leq\biggl\{\sum_{k=1}^{K}\max\bigl\{\bigl(\sum_{i\in\{i\mid\delta_{ik}>0\}}\delta_{ik}\psi_{H}(G_{i})\bigr)^{2},\ \bigl(\sum_{i\in\{i\mid\delta_{ik}<0\}}\delta_{ik}\psi_{H}(G_{i})\bigr)^{2}\bigr\}\biggr\}^{1/2}\eqqcolon\mathrm{UB}(H).$ (9)
We here only describe the sketch of the proof (the proof is in Appendix B). An essential idea is that we expand $\boldsymbol{g}_{H}$ as a linear combination of the derivatives with respect to the GML intermediate representation $h_{k}$, i.e., $\delta_{ik}$, by which the $k$-th element of the gradient can be written as $(\boldsymbol{g}_{H})_{k}=\sum_{i=1}^{n}\delta_{ik}\psi_{H}(G_{i})$. By using the monotonically non-increasing property of $\psi_{H}(G_{i})$, i.e., $\psi_{H^{\prime}}(G_{i})\leq\psi_{H}(G_{i})$ if $H\sqsubseteq H^{\prime}$, we can derive ${\rm UB}(H)$. Note that, usually, there is no need to explicitly consider this specific expansion of $(\boldsymbol{g}_{H})_{k}$; the theorem can only be derived by deliberately reducing it to the linear combination of $\delta_{ik}$.
This theorem indicates that, for any $H^{\prime}$ that contains $H$ as a subgraph, the L2 norm of the gradient $\|\boldsymbol{g}_{H^{\prime}}\|_{2}$ can be bounded by $\mathrm{UB}(H)$. Note that the upper bound $\mathrm{UB}(H)$ can be calculated without generating $H^{\prime}$. From the rule (7) and Theorem 2.1, we can immediately obtain the following important rule:
@@
For $\forall H^{\prime}\in\{H^{\prime}\mid H^{\prime}\sqsupseteq H,H^{\prime}\in\overline{{\cal W}}\}$,
$\mathrm{UB}(H)\leq\lambda\text{ and }H\in\overline{{\cal W}}\ \ \Rightarrow\ \ \mathrm{prox}\left(\boldsymbol{\beta}_{H^{\prime}}-\eta\ \boldsymbol{g}_{H^{\prime}}\right)=\boldsymbol{0}$ (10)
This is a direct consequence from the fact that if the conditions in the left side of (10) holds, we have $\|\boldsymbol{g}_{H^{\prime}}\|_{2}\leq\mathrm{UB}(H)\leq\lambda$ from (9). As a result, by using (7), we obtain (10). Corollary (2.1) means that if the condition in (10) holds, for $\forall H^{\prime}$ that contains $H$ as a subgraph, we can omit the update. Further, another notable advantage of this rule is that it does not depend on the step size $\eta$ (even when an iterative backtrack algorithm is employed for $\eta$, the evaluation of (10) is not required to repeat for every backtrack iteration).
1function Train-EIN($\Lambda$)

### v2

# Exact Subgraph Isomorphism Network with Mixed $L_{0,2}$ Norm Constraint for Predictive Graph Mining

@@ Abstract
In the graph-level prediction task (predict a label for a given graph), the information contained in subgraphs of the input graph plays a key role. In this paper, we propose Exact subgraph Isomorphism Network (EIN), which combines the exact subgraph enumeration, a neural network, and a sparse regularization by the mixed $L_{0,2}$ norm constraint. In general, building a graph-level prediction model achieving high discriminative ability along with interpretability is still a challenging problem. Our combination of the subgraph enumeration and neural network contributes to high discriminative ability about the subgraph structure of the input graph. Further, the sparse regularization in EIN enables us 1) to derive an effective pruning strategy that mitigates computational difficulty of the enumeration while maintaining the prediction performance, and 2) to identify important subgraphs that contributes to high interpretability. We empirically show that EIN has sufficiently high prediction performance compared with standard graph neural network models, and also, we show examples of post-hoc analysis based on the selected subgraphs.
@@ 1 Introduction
EIN consists of Graph Mining Layer (GML) and Feed Forward Network (FFN). GML implements a mechanism that identifies only $s$ subgraphs important for the prediction. This is achieved through a group sparse constraint (Yuan and Lin,, 2006), where $s$ is a pre-specified small integer (typically, less than $100$). We regard a set of neural network weights from an input unit (corresponding to one subgraph) to the next layer as a group, by which the adaptive subgraph selection can be realized. Only the selected subgraphs are used in the subsequent FFN that learns the dependency between the selected subgraphs and the prediction target. Although the number of subgraphs becomes small in the trained model, this approach still suffers from the computational difficulty because all the candidate subgraphs are still required to consider during the optimization process.
To tackle the computational difficulty, we combine a subgraph enumeration by graph mining (Yan and Han,, 2002) and the iterative hard thresholding (IHT) (e.g., Blumensath and Davies,, 2008; Herrity et al.,, 2006; Jain et al.,, 2014; Damadi et al.,, 2024), by which a pruning strategy for unnecessarily subgraphs can be derived, as illustrated in 1 (c). In IHT, only when the norm of the gradient corresponding to a subgraph $H$ ranks within the top $s$-th largest values among all the possible subgraphs, parameters for that subgraph is required to update. We show that, by deriving an upper bound of the norm of the gradient, an efficient pruning strategy for the subgraph enumeration can be constructed. Our pruning strategy has the following two important implications. First, this pruning enables us to train EIN without enumerating all the candidate subgraphs, which makes EIN computationally tractable. Second, our pruning strategy maintains the quality of the prediction compared with when we do not perform the pruning. This is because we only omit the computations that does not have any effect on the prediction.
@@
  • We propose EIN, which is a neural network model that uses the exact subgraph isomorphism feature. Through a group sparse constraint, we formulate EIN so that a small number of subgraphs can be identified by which an insight about the important substructure can be extracted for the given graph-level prediction task.
  • We show that the combination of the subgraph enumeration and IHT with backpropagation can derive an efficient pruning strategy that makes the EIN training computationally tractable. We further reveal that our pruning strategy does not degrade the prediction quality, and the resulting algorithm has at least sublinear convergence rate to a critical point.
@@ 2 Proposed Method: Exact Subgraph Isomorphism Network
Our proposed method, called Exact subgraph Isomorphism Network (EIN), considers the classification problem of a graph $G\in{\mathcal{G}}$, where $\mathcal{G}$} is a set of labeled graphs. $G$ consists of a set of nodes and edges between nodes and each node can have a categorical label. The training data is $\{(G_{i},y_{i})\}_{i\in[n]}$, where $y_{i}\in{\mathcal{Y}}$ is a graph label, $n$ is the number of instances, and $n]=\{1,\ldots,n\}$. Here, although we only focus on the classification problem, the regression problem can also be handled by just replacing the loss function.
@@ 2.1 Model Definition
As shown in Fig. 1 (b), EIN consists of Graph Mining Layer (GML) and Feed Forward Network (FFN). Let $K$ be the number of output units of GML. We define $\mathrm{GML}:{\mathcal{G}}\rightarrow\mathbb{R}^{K}$ as follows.
@@
$\boldsymbol{h}$ $=\sum_{H\in{\mathcal{H}}}\boldsymbol{\beta}_{H}\psi_{H}(G_{i})+\boldsymbol{b},$
where $\boldsymbol{B}\in\mathbb{R}^{K\times|{\mathcal{H}}|}$ is a weight matrix whose columns consist of $\boldsymbol{\beta}_{H}\in\mathbb{R}^{K}$ for $H\in{\mathcal{H}}$, $\boldsymbol{b}\in\mathbb{R}^{K}$ is a bias parameter, and $\sigma:\mathbb{R}^{K}\rightarrow\mathbb{R}^{K}$ is an activation function (this time, we use the standard sigmoid function). Each $\boldsymbol{\beta}_{H}$ can be seen as a representation corresponding to each subgraph $H$. The entire prediction model is defined as
@@
where $\boldsymbol{\theta}$ is parameters of FFN. We use the cross entropy loss function by which our objective function is defined as $\ell(\boldsymbol{B},\boldsymbol{b},\boldsymbol{\theta})\coloneqq\sum_{i=1}^{n}\mathrm{loss}(y_{i},f(G_{i});\boldsymbol{B},\boldsymbol{b},\boldsymbol{\theta})$. We optimize the parameters $\boldsymbol{B},\boldsymbol{b}$, and $\boldsymbol{\theta}$ through the following optimization problem in which a group sparsity constraint is imposed on $\boldsymbol{B}$:
$\begin{split}\min_{\boldsymbol{B},\boldsymbol{b},\boldsymbol{\theta}}\ &\ell(\boldsymbol{B},\boldsymbol{b},\boldsymbol{\theta})\\ \text{s.t.}\ &\|\boldsymbol{B}\|_{0,2}\leq s,\end{split}$ (2)
where
$\|\boldsymbol{B}\|_{0,2}=|\{H\mid\|\boldsymbol{\beta}_{H}\|_{2}\neq 0\}|$
is the mixed $L_{0,2}$ norm, and $s>0$ is the constant parameter. The constraint in (2) indicates that $\boldsymbol{B}$ only has $s$ non-zero columns. In other words, this means that only $s$ subgraphs are used in $f$, by which we can identify important predictive subgraphs as $\{H\mid\boldsymbol{\beta}_{H}\neq\boldsymbol{0}\}$. However, since the size of $\boldsymbol{B}$, i.e., $K\times|{\mathcal{H}}|$, is quite large, naïve optimization of (2) can be difficult.
@@ 2.2 Optimization
Our optimization algorithm is based on the block coordinate descent, in which each one of $\boldsymbol{B},\boldsymbol{b}$, and $\boldsymbol{\theta}$ are updated alternately while the other two parameters are fixed. Since $\boldsymbol{B}$ has the group sparse constraint, we apply iterative hard thresholding (IHT) to $\boldsymbol{B}$ (e.g., Beck and Eldar,, 2013; Bao et al.,, 2014; Yuan et al.,, 2018; Damadi and Shen,, 2022). On the other hand, $\boldsymbol{b}$ and $\boldsymbol{\theta}$ have no sparse penalty and we can simply apply the standard gradient descent.
We iterate the following procedure, in which $t$ is the number of iterations:
  1. 1.
    Update $\boldsymbol{B}$ through the hard thresholding (HT), defined as
    $\begin{split}\boldsymbol{B}^{t+\frac{1}{2}}&=\boldsymbol{B}^{t}-\gamma^{t}_{\boldsymbol{B}}\nabla_{\boldsymbol{B}}\ell^{t}\\ \boldsymbol{B}^{t+1}&=H_{s}(\boldsymbol{B}^{t+\frac{1}{2}}),\end{split}$ (3)
    where $\nabla_{\boldsymbol{B}}\ell^{t}={\frac{\partial\ell(\boldsymbol{B}^{t},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})}{\partial\boldsymbol{B}}}$ is the gradient of $\boldsymbol{B}$, $\gamma^{t}_{\boldsymbol{B}}$ is a step length, and
    $H_{s}(\boldsymbol{B})=\mathop{\text{argmin}}_{\|\boldsymbol{Z}\|_{0,2}\leq s}\|\boldsymbol{Z}-\boldsymbol{B}\|_{F}$
    is the HT operator. The step length parameter $\gamma^{t}_{\boldsymbol{B}}$ is determined by the backtrack search (Nocedal and Wright,, 2006).
  2. 2.
    Update $\boldsymbol{b}^{t}$ and $\boldsymbol{\theta}^{t}$ to $\boldsymbol{b}^{t+1}$ and $\boldsymbol{\theta}^{t+1}$, respectively, by using the standard gradient descent.
Since $\mathcal{H}$} contains a large number of subgraphs, calculating (3) for the entire $\boldsymbol{B}\in\mathbb{R}^{K\times|{\mathcal{H}}|}$ is not directly tractable. In § 2.2.1, we derive a pruning rule by which we can perform (3) without enumerating all the subgraphs in $\mathcal{H}$}. § 2.2.2 describes the entire procedure, and § 2.2.3 provides a theoretical convergence analysis.
@@ 2.2.1 Backpropagation with Gradient Pruning
By combining a graph mining algorithm and the HT operation, we can build an efficient pruning algorithm. We employ a well-known graph mining algorithm called gSpan (Yan and Han,, 2002) that can efficiently enumerate all the subgraphs in a given set of graphs. Figure 1 (c) is an illustration of a graph mining tree. At each tree node, gSpan expands the graph (add an edge and a node) as far as the expanded graph is included in the given dataset as a subgraph. By providing a unique code to each generated graphs (called the DFS code because it is based on depth-first search in a graph), gSpan can enumerate all the subgraphs without generating duplicated graphs (For further detail, see (Yan and Han,, 2002)). As shown in Fig. 1 (c), all the graphs in the tree contains their ancestors as a subgraph.
It is easy to see that the HT operator simply selects the columns with the top-$s$ largest norm because
$H_{s}(\boldsymbol{B})$ $=\mathop{\text{argmin}}_{\|\boldsymbol{Z}\|_{2,0}\leq s}\|\boldsymbol{Z}-\boldsymbol{B}\|_{F}^{2}=\mathop{\text{argmin}}_{\|\boldsymbol{Z}\|_{2,0}\leq s}\sum_{k=1}^{|{\mathcal{H}}|}\|\boldsymbol{Z}_{k}-\boldsymbol{B}_{k}\|_{2}^{2},$
where $\boldsymbol{B}_{k}$ and $\boldsymbol{Z}_{k}$ is the $k$-th column of $\boldsymbol{B}$ and $\boldsymbol{Z}$, respectively. Let $H_{1},\ldots,H_{|{\mathcal{H}}|}$ be the subgraphs such that $\|\boldsymbol{\beta}^{t+\frac{1}{2}}_{H_{1}}\|_{2}>\|\boldsymbol{\beta}^{t+\frac{1}{2}}_{H_{2}}\|_{2}>\cdots>\|\boldsymbol{\beta}^{t+\frac{1}{2}}_{H_{|{\mathcal{H}}|}}\|_{2}$, where $\boldsymbol{\beta}^{t+\frac{1}{2}}_{H}$ is a column of $\boldsymbol{B}^{t+\frac{1}{2}}$. Then, $\boldsymbol{\beta}^{t+1}_{H}$, which is a column of $\boldsymbol{B}^{t+1}$, is written as
$\boldsymbol{\beta}^{t+1}_{H}=\begin{cases}\boldsymbol{\beta}^{t+\frac{1}{2}}_{H}&\text{ if }H\in\{H_{1},\ldots,H_{s}\},\\ \boldsymbol{0}&\text{ otherwise. }\end{cases}$
This means that we only need to identify $H$ with the top-$s$ largest $\|\boldsymbol{\beta}^{t+\frac{1}{2}}_{H}\|_{2}$ during the graph mining tree traverse because $\boldsymbol{\beta}^{t+1}_{H}$ for the other $H$ is $\boldsymbol{0}$.
Let $S_{t}$ be the set of selected $H$ at the iteration $t$. Since $|S_{t}|=s$, which is assumed to be a small value (in our later experiments, $s\leq 100$), it is not computationally difficult to calculate $\boldsymbol{\beta}^{t+\frac{1}{2}}_{H}$ for $H\in S_{t}$. On the other hand, we have an intractable number of $H\notin S_{t}$. Therefore, we remove $H$ that does not have a possibility of being included in $S_{t+1}$. Since $\boldsymbol{\beta}^{t}_{H}=\boldsymbol{0}$ for $H\notin S_{t}$, if $H\notin S_{t}$, we see $\|\boldsymbol{\beta}_{H}^{t+\frac{1}{2}}\|_{2}=\gamma_{\boldsymbol{B}}^{t}\|\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}\|_{2}$, where $\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}={\frac{\partial\ell(\boldsymbol{B}^{t},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})}{\partial\boldsymbol{\beta}_{H}}}$. To define our pruning strategy, the following upper bound of $\|\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}\|_{2}$ plays a key role:
@@
Let $H^{\prime}\sqsupseteq H$ and $\delta^{t}_{ik}={\frac{\partial\mathrm{loss}(y_{i},f(G_{i});\boldsymbol{B}^{t},\boldsymbol{b}^{t},\boldsymbol{\theta}^{t})}{\partial h_{k}}}$. Then,
$\|\nabla_{\boldsymbol{\beta}_{H^{\prime}}}\ell^{t}\|_{2}\leq$ $\biggl\{\sum_{k=1}^{K}\max\bigl\{\bigl(\sum_{i\in\{i\mid\delta^{t}_{ik}>0\}}\delta^{t}_{ik}\psi_{H}(G_{i})\bigr)^{2},$
$\ \ \bigl(\sum_{i\in\{i\mid\delta^{t}_{ik}<0\}}\delta^{t}_{ik}\psi_{H}(G_{i})\bigr)^{2}\bigr\}\biggr\}^{1/2}\eqqcolon\mathrm{UB}(H).$ (4)
We here only describe the sketch of the proof (the proof is in Appendix C). An essential idea is that we expand $\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}$ as a linear combination of the derivatives with respect to the GML intermediate representation $h_{k}$, i.e., $\delta^{t}_{ik}$, by which the $k$-th element of the gradient can be written as $(\nabla_{\boldsymbol{\beta}_{H}}\ell^{t})_{k}=\sum_{i=1}^{n}\delta^{t}_{ik}\psi_{H}(G_{i})$. By using the monotonically non-increasing property of $\psi_{H}(G_{i})$, i.e., $\psi_{H^{\prime}}(G_{i})\leq\psi_{H}(G_{i})$ if $H\sqsubseteq H^{\prime}$, we can derive ${\rm UB}(H)$. Note that, usually, there is no need to explicitly consider this specific expansion of $(\nabla_{\boldsymbol{\beta}_{H}}\ell^{t})_{k}$; the theorem can only be derived by deliberately reducing it to the linear combination of $\delta^{t}_{ik}$.
Theorem 2.1 indicates that, for any $H^{\prime}$ that contains $H$ as a subgraph, the $L_{2}$ norm of the gradient $\|\nabla_{\boldsymbol{\beta}_{H^{\prime}}}\ell^{t}\|_{2}$ can be bounded by $\mathrm{UB}(H)$. Note that the upper bound $\mathrm{UB}(H)$ can be calculated without generating $H^{\prime}$. From Theorem 2.1, we can immediately obtain the following pruning rule with a threshold parameter $\eta$:
@@
Let $\zeta\leq\|\boldsymbol{\beta}^{t+\frac{1}{2}}_{H_{s}}\|_{2}/\gamma_{\boldsymbol{B}}^{t}$, meaning that $\zeta$ is less than or equal to the $s$-th largest value of $\|\boldsymbol{\beta}^{t+\frac{1}{2}}_{H}\|_{2}/\gamma_{\boldsymbol{B}}^{t}$ among $H\in{\mathcal{H}}$. For $\forall H^{\prime}\in\{H^{\prime}\mid H^{\prime}\sqsupseteq H,\ H^{\prime}\notin S_{t}\}$,
$\mathrm{UB}(H)<\zeta\ \ \Rightarrow\ \ \boldsymbol{\beta}_{H^{\prime}}^{t+1}=\boldsymbol{0}$ (5)
This is a direct consequence from the fact that if the conditions in (5) holds, we have
$\|\boldsymbol{\beta}^{t+\frac{1}{2}}_{H^{\prime}}\|_{2}=\gamma_{\boldsymbol{B}}^{t}\|\nabla_{\boldsymbol{\beta}_{H^{\prime}}}\ell^{t}\|_{2}\leq\gamma_{\boldsymbol{B}}^{t}\mathrm{UB}(H)<\gamma_{\boldsymbol{B}}^{t}\zeta\leq\|\boldsymbol{\beta}^{t+\frac{1}{2}}_{H_{s}}\|_{2}$
from (4). As a result, we see that $\|\boldsymbol{\beta}^{t+\frac{1}{2}}_{H^{\prime}}\|_{2}$ cannot be ranked in the top-$s$ largest values. Therefore, we can omit the update for $\forall H^{\prime}$ that contains $H$ as a subgraph if $\mathrm{UB}(H)<\zeta$ holds.
To use the pruning rule Corollary 2.1, we need to determine $\zeta$. We first consider the case that the step length $\gamma_{\boldsymbol{B}}^{t}$ is fixed beforehand, in which a simple approach can be applied. First, $\|\boldsymbol{\beta}_{H}^{t+\frac{1}{2}}\|$ is calculated for all $H\in S_{t}$. After that, the graph mining tree is traversed during which each one of $\|\boldsymbol{\beta}_{H}^{t+\frac{1}{2}}\|$ is calculated for the subgraph $H$ of each visited tree node. Suppose that ${\mathcal{H}}_{\rm visited}$ represents the union of $S_{t}$ and a set of $H$ corresponding to the already visited tree nodes. The threshold of the pruning rule (5) can be set as $\zeta=s\operatorname{-max}\{\|\boldsymbol{\beta}_{H}^{t+\frac{1}{2}}\|/\gamma_{\boldsymbol{B}}^{t}\}_{H\in{\mathcal{H}}_{\rm visited}}$, where $s\operatorname{-max}$ represents the $s$-th largest value, by which we can guarantee $\zeta\leq\|\boldsymbol{\beta}^{t+\frac{1}{2}}_{H_{s}}\|_{2}/\gamma_{\boldsymbol{B}}^{t}$. As a result, if $\mathrm{UB}(H)<\zeta$ holds for the threshold $\zeta$ at any node during the tree traverse, we can discard the entire subtree below the current node. However, when we use a backtrack strategy for the step length, we cannot calculate $\|\boldsymbol{\beta}_{H}^{t+\frac{1}{2}}\|_{2}$ before the update of $\boldsymbol{B}^{t}$ because $\gamma_{\boldsymbol{B}}^{t}$ is not determined yet.
Instead, we can also build a threshold $\zeta$ without using $\gamma_{\boldsymbol{B}}^{t}$ as follows.
$s\operatorname{-max}\{\ \|\boldsymbol{\beta}_{H}^{t+\frac{1}{2}}\|_{2}/\gamma_{\boldsymbol{B}}^{t}\ \}_{H\in{\mathcal{H}}_{\rm visited}}$
$=s\operatorname{-max}\left\{\{\frac{1}{\gamma_{\boldsymbol{B}}^{t}}\|\boldsymbol{\beta}_{H}^{t}-\gamma_{\boldsymbol{B}}^{t}\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}\|_{2}\}_{H\in S_{t}}\cup\{\|\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}\|_{2}\}_{H\in{\mathcal{H}}_{\rm visited}\setminus S_{t}}\right\}$
$\geq s\operatorname{-max}\left\{\eta_{H}\right\}_{H\in{\mathcal{H}}_{\rm visited}},$ (6)
where
$\eta_{H}\coloneqq\begin{cases}\min_{\gamma\in\Gamma}\frac{1}{\gamma}\|\boldsymbol{\beta}_{H}^{t}-\gamma\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}\|_{2}&\text{ if }H\in S_{t},\\ \|\nabla_{\boldsymbol{\beta}_{H}}\ell^{t}\|_{2}&\text{ if }H\notin S_{t},\end{cases}$
and $\Gamma$ is a candidate space of the step length $\gamma_{\boldsymbol{B}}^{t}$. We use this lower bound (6) as $\zeta$. In the derivation of (6), $H\in S_{t}$ and other $H\in{\mathcal{H}}_{\rm visited}\setminus S_{t}$ are separately considered. For $H\in{\mathcal{H}}_{\rm visited}\setminus S_{t}$ (i.e., $H\notin S_{t}$), $\gamma_{\boldsymbol{B}}^{t}$ is vanished in $\eta_{H}$, and therefore, the calculation does not depend on the step length. On the other hand, for $H\in S_{t}$, since the step length cannot be directly removed, we derive a lower bound with respect to possible step lengths. This lower bound $\min_{\gamma\in\Gamma}$ can be easily calculated (See Appendix D for detail).
1function Train-EIN($\mathcal{S}$})
2 $\boldsymbol{B}\leftarrow\boldsymbol{0}$
3   Initialize $\boldsymbol{b}$ and $\boldsymbol{\theta}$ randomly